OpenStudy (anonymous):

Ok, if i dont go from 13.4 percent to 60 percent by thursday, i fail, will someone help me please!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

good gosh

OpenStudy (anonymous):

*puts in ammo* BOOM!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

for what class is this flvs

OpenStudy (anonymous):

no, edgenuity.com

OpenStudy (anonymous):

how did you even get a 13.4, what math are you in

OpenStudy (anonymous):

no, im 13.4 percent done with the course, my grade is a b

OpenStudy (anonymous):

oh alright i was about to say, so what do you need help in

OpenStudy (anonymous):

math

OpenStudy (anonymous):

what math are you taking

OpenStudy (anonymous):

???

OpenStudy (anonymous):

hard math?

OpenStudy (anonymous):

what grade are you in

OpenStudy (anonymous):

6th, this isnt the first time i failed

OpenStudy (anonymous):

see no-one really cares the majority arlready left

OpenStudy (anonymous):

im wasting time right now

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

kill me

OpenStudy (anonymous):

IS ANYONE EVEN LISTENING!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

you even haven't gave us a question to help you on

OpenStudy (anonymous):

and if you have equations and stuff like that go to mathway.com that helps me alot

OpenStudy (anonymous):

IVE TRIED IT IT SUCKS

OpenStudy (anonymous):

KILL ME ALREADY

OpenStudy (anonymous):

well post questions and we might be able to help you, its been awhile since 6th grade math but i can try

OpenStudy (anonymous):

Mrs. Perez is buying a refrigerator for $1368.25, which includes taxes and delivery charges. She pays $497.80 in cash and will write a check for the balance. For how much should she write the check? $870.35 $870.45 $971.45 $1866.05

OpenStudy (anonymous):

I told you that answer its b

OpenStudy (anonymous):

my comp froze earlier so idint get it

OpenStudy (anonymous):

thank you...

OpenStudy (anonymous):

all you do is divide $1368.25 by $497.80 and get b your welcom

OpenStudy (anonymous):

here is another one

OpenStudy (anonymous):

Which number is composite? 19 23 33 37

OpenStudy (anonymous):

b

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Carlotta drove 72.6 miles on Monday. She drove 84.18 miles on Tuesday. Which gives the best estimate for how many more miles Carlotta drove on Tuesday? 84 – 73 = 11 miles 73 – 84 = 11 miles 85 – 72 = 13 miles 84 – 71 = 13 miles

OpenStudy (anonymous):

b

OpenStudy (anonymous):

is b the answer for all of them/? lol

OpenStudy (anonymous):

yes need explanation for the last one

OpenStudy (anonymous):

its weird that all of them so far are b, lol

OpenStudy (anonymous):

Harshini rode her bicycle 13.76 kilometers on Monday, 9.8 kilometers on Wednesday, and 16.415 kilometers on Friday. She used the addition below to determine the total distance she rode over the three days.

OpenStudy (anonymous):

you round 72.6 to 73 and 84.16 stays 84, and so 84-73=11

OpenStudy (anonymous):

where is the rest of the question

OpenStudy (anonymous):

Which method determines whether her answer is reasonable? Remove the decimal points and add 137 + 98 + 16415 = 17,889; since 17,889 is not close to 39.975, her answer is not reasonable. Round all the numbers to 20; since 20 + 20 + 20 = 60 and 60 is not close to 39.975, her answer is not reasonable. Round all the numbers to 10; since 10 + 10 + 10 = 30 and 30 is close to 39.975, her answer is reasonable. Round 13.76 to 14, round 9.8 to 10, and round 16.415 to 16; since 14 + 10 + 16 = 40 and 40 is close to 39.975, her answer is reasonable.

OpenStudy (anonymous):

@hershey21

OpenStudy (anonymous):

the last one sorry

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Which number is prime? 49 51 53 55

OpenStudy (anonymous):

53

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

What is the prime factorization of 84?

OpenStudy (anonymous):

Which number is composite? 19 23<------its no this one sorry made a mistake on this one only 33 37<-------answer

OpenStudy (anonymous):

those are the answer choices

OpenStudy (anonymous):

its 2^2x3x7

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

heres another

OpenStudy (anonymous):

Kia thinks that every number has at least two factors. Which explains whether Kia is correct? Kia is correct because prime numbers have two factors and composite numbers have more than two factors. Kia is correct because composite numbers have two factors and prime numbers have more than two factors. Kia is incorrect because prime numbers do not have any factors. Kia is incorrect because the only factor of 1 is 1.

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

I have to go sorry

OpenStudy (anonymous):

thsi was wrong

OpenStudy (anonymous):

Mr. Atlas, the weatherman, reports that the temperature at 6 a.m. was 44.65 degrees and the temperature at 10 a.m. is 52.14 degrees. Which is the best estimate of the number of degrees that the temperature has risen? 7 7.49 8 8.49

OpenStudy (anonymous):

u said d and it was a

OpenStudy (anonymous):

this one too

OpenStudy (anonymous):

Which number is composite? 19 23 33 37

OpenStudy (anonymous):

is was c not d

OpenStudy (anonymous):

plus this

OpenStudy (anonymous):

Carlotta drove 72.6 miles on Monday. She drove 84.18 miles on Tuesday. Which gives the best estimate for how many more miles Carlotta drove on Tuesday? 84 – 73 = 11 miles 73 – 84 = 11 miles 85 – 72 = 13 miles 84 – 71 = 13 miles

OpenStudy (anonymous):

@marihelenh you there?

OpenStudy (anonymous):

that last one was a not b

OpenStudy (marihelenh):

What? Oh sorry, I was busy with other stuff. What do you need help with?

OpenStudy (anonymous):

and this one was d not a

OpenStudy (anonymous):

Kia thinks that every number has at least two factors. Which explains whether Kia is correct? Kia is correct because prime numbers have two factors and composite numbers have more than two factors. Kia is correct because composite numbers have two factors and prime numbers have more than two factors. Kia is incorrect because prime numbers do not have any factors. Kia is incorrect because the only factor of 1 is 1.

OpenStudy (marihelenh):

What question do you need help with?

OpenStudy (anonymous):

well i already faled the test review, goo thing it didnt count

OpenStudy (marihelenh):

Oh jeez, I'm sorry to hear that. Were these all questions on it that other people helped you with?

OpenStudy (anonymous):

yes, five of them were incorrect

OpenStudy (anonymous):

out of twelve

OpenStudy (anonymous):

kill me, please @marihelenh

OpenStudy (marihelenh):

Oh no, no, no. There is no need for that. I'm looking through these and I feel like I would have gotten them correct.

OpenStudy (anonymous):

...

OpenStudy (anonymous):

-.-, no-one really cares about a teenagers problems

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

ive basiclly already failed, why do i bother

OpenStudy (anonymous):

@marihelenh

OpenStudy (marihelenh):

Teenager problems? I happen to still be a teen, and I know what you're going through, although I didn't have Open Study to help me. I went through the previous questions and who was ever helping had no clue what they were doing. Of the questions you had on here, I already found at least 4 or 5 incorrect answers.

OpenStudy (anonymous):

5

OpenStudy (marihelenh):

Is there any questions you need me to help you with?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok, this is a topic test, so i have 2 hours to finish 25 problems

OpenStudy (marihelenh):

First, what topic? And is it a pdf or anything?

OpenStudy (anonymous):

idk its math, like prime numbers and the stuff i just asked

OpenStudy (marihelenh):

Can we start it in like 5 minutes? That way I can get my stuff handled so I don't break right in the middle?

OpenStudy (anonymous):

ok

OpenStudy (marihelenh):

Alright, let's do this.

OpenStudy (anonymous):

ok...